Solve the system by the substitution method.
3x/2+y/5=-3
4x+y/3=-11

Answer by stanbon(75887)

3x/2+y/5=-3
4x+y/3=-11
------------------
15x + 2y = -30
12x + y = -33
----------------
15x + 2y = -30
y = -12x - 33
-------
Substitute for "y" and solve for "x":
15x + 2(-12x-33) = -30
15x -24x-66 = -30
-9x = 36
x = -4
-------
Solve for "y":
y = -12x-33
y = -12*-4 - 33
y = 48 - 33
y = 15
=================
Cheers,
Stan H.